OpenStudy (anonymous):

for a certain computer desk, the manufacturing cost C per desk ( in dollars) is C= 500x+ 5000 over x where x is the number of desks manufactured, find the average cost per desk when manufacturing 200 desks

OpenStudy (amistre64):

\[C(x) = 500x +\frac{5000}{x}\] is this the equation?

OpenStudy (anonymous):

no, C = 500x + 5000 _______________ X

OpenStudy (amistre64):

so we simply plug in 200 for x: 500(200) + 5000 = 100000 + 5000 = 105000 105000/2 = 52500

OpenStudy (amistre64):

opps; /200 lol

OpenStudy (amistre64):

105000/200 = 1050/2 = 525

OpenStudy (anonymous):

dont understand

OpenStudy (anonymous):

dont get 525

OpenStudy (amistre64):

when x = 200: 500x + 5000 ------------ x 500(200) + 5000 100000 + 5000 105000 --------------- = --------------- = ------- 200 200 200 105000 1050 ------- = ------ = 525 200 2